OpenStudy (compassionate):

Let f(x) = 2x - 6. Solve f-1(x) when x = 2. Okay... so we need the inverse of x. So f(x) can just be set to y... right? y = 2x - 6 2x = y - 6 x = y - 3 2 = y - 3 y = -5? Or is this completely wrong?

OpenStudy (rational):

y = 2x - 6 ` 2x = y - 6` mistake x = y - 3 2 = y - 3 y = -5?

OpenStudy (compassionate):

-2x = -y - 6?

OpenStudy (compassionate):

Because I added the first time, so would I need to subtract instead?

OpenStudy (rational):

yes y = 2x - 6 to isolate x, first step is to add 6 both sides : y + 6 = 2x next, divide 2 both sides : (y+6)/2 = x

OpenStudy (rational):

now you can switch x and y : (x+6)/2 = y

OpenStudy (rational):

so the inverse function is : \[\large y = (x+6)/2\]

OpenStudy (rational):

plugin x = 2

OpenStudy (compassionate):

So then 2 + 6/2 which is just 8/2 = 4

OpenStudy (compassionate):

So lets do a practice one to make sure I got it right Let f(x) = 2x + 2. Solve f-1(x) when x = 4. y = 2x + 2 y - 2 = 2x (y - 2/2) = x Plug in 4. 4 - 2 /2 = x x = 1

OpenStudy (rational):

thats right ! but don't forget about parenthesis

OpenStudy (rational):

Let f(x) = 2x + 2. Solve f-1(x) when x = 4. y = 2x + 2 y - 2 = 2x `(y - 2/2) = x` your teacher cries at this step Plug in 4. 4 - 2 /2 = x x = 1

OpenStudy (rational):

Corrected below : Let f(x) = 2x + 2. Solve f-1(x) when x = 4. y = 2x + 2 y - 2 = 2x `(y - 2)/2 = x` Plug in 4. `(4 - 2) /2 = x` x = 1

OpenStudy (compassionate):

Techy wechy. Thanks!

OpenStudy (rational):

yw!